OpenStudy (anonymous):

If r=6cm and h=2cm, give or take 0.1cm, which is the expected maximum percentage error in calculating V= pie*(r^2)*h

OpenStudy (anonymous):

Apply the propagation of error to the equation of volume. Are you familiar with it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

could you show me?

OpenStudy (anonymous):

Sure. Actually, for worst case, its a complete differential. My mistake. But here we go. Percent errors come in the form of dr/r, which in our case is .1/6, and dh/h is .1/2. We are looking for dV/V. Do you know any easy way to get to this result?

OpenStudy (anonymous):

no, could you go through it quickly all the way to the end please?

OpenStudy (anonymous):

Sure. Take the natural log of both sides, giving you: \[lnV = \ln(\pi r ^{2} h) = \ln \pi + 2 \ln r + \ln h\] Can you take the derivative of this?

OpenStudy (anonymous):

(1/pie)+(2/r)+(1/h)?

OpenStudy (anonymous):

Close. Each variable is subject to change, however.

OpenStudy (anonymous):

oh wait is it 2/r?

OpenStudy (anonymous):

\[dV/V = 2dr/r + dh/h\] ln pi is a number, so it falls away when a derivative is taken.

OpenStudy (anonymous):

Since we are looking for the worst case, you take the absolute value of each term. This doesn't effect this set up at all, however. So we know dr/r and dh/h. So now we plug in and get our answer for dV/V.

OpenStudy (anonymous):

so answer is 6%?

OpenStudy (anonymous):

2*(.1/6) + (.1/2) = .083333 = 8.33%

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

Anytime. Take care!